I am having a really hard time trying to learn Algebra. Need assistace Please.
For the first phase of the project, imagine you want to cover the backyard with decorative rock and plant some trees. You need 30 tons of rock to cover the area. If each ton costs $60 and each tree is $84, what is the maximum number of trees you can buy with a budget of $2,500? Write an inequality that illustrates the problem and solve.

Answer by stanbon(57246)

For the first phase of the project, imagine you want to cover the backyard with decorative rock and plant some trees. You need 30 tons of rock to cover the area. If each ton costs $60 and each tree is $84, what is the maximum number of trees you can buy with a budget of $2,500? Write an inequality that illustrates the problem and solve.
----------------------
Money remaining after rock is purchased:
2500 - 60*30 = $700
-----
Number of trees you can purchase = 700/84 = 8 1/3
Round down to 8 trees.
------
Inequality
# of trees <= [2500-60*30]/84
==================================
Cheers,
Stan H.